# Fair and Efficient Completion of Indivisible Goods

### v1

@@ Abstract
We formulate the problem of fair and efficient completion of indivisible goods, defined as follows: Given a partial allocation of indivisible goods among agents, does there exist an allocation of the remaining goods (i.e., a completion) that satisfies fairness and economic efficiency guarantees of interest? We study the computational complexity of the completion problem for prominent fairness and efficiency notions such as envy-freeness up one good (EF1111), proportionality up to one good (Prop1111), maximin share (MMS), and Pareto optimality (PO), and focus on the class of additive valuations as well as its subclasses such as binary additive and lexicographic valuations. We find that while the completion problem is significantly harder than the standard fair division problem (wherein the initial partial allocation is empty), the consideration of restricted preferences facilitates positive algorithmic results for threshold-based fairness notions (Prop1111 and MMS). On the other hand, the completion problem remains computationally intractable for envy-based notions such as EF1111 and EF1111+PO even under restricted preferences.
@@ 1 Introduction
[…]
Table 1: Summary of our results for indivisible goods. Each row corresponds to a fairness/efficiency property. Each column corresponds to a preference restriction and whether we are looking at the standard or the completion problem; ‘standard’ is a special case of completion problem when there are no frozen goods. Each cell shows the computational complexity of the standard or completion problem under specific scenarios.
@@
  • For general additive valuations, we show that the completion problem for EF1111 and Prop1111 (as well as their combinations with Pareto optimality) is NP-hard. This hardness result holds even for two agents with additive valuations and for three agents with identical additive valuations. We also establish the hardness of deciding the existence of an MMS allocation, a problem whose lower bound on complexity is open in the standard setting. Moreover, we demonstrate that $\alpha$-MMS may not exist for every $\alpha>0$ even when the initial allocation is Pareto optimal.
@@ 3 Binary Additive Valuations
In the standard setting of fair division (i.e., without any frozen goods), an EF1111 allocation always exists and can be efficiently computed by a simple round-robin algorithm [CKM+19unreasonable]. However, as discussed in Section 1, an EF1111 allocation could fail to exist when some goods are pre-allocated. Our first main result in this section is that EF1111-Completion is computationally intractable even when the agents have binary additive valuations and even when the frozen allocation is Pareto optimal. Note that, for binary additive valuations, an allocation is Pareto optimal if and only if each good is allocated to an agent who approves it.
@@
If there is no frozen good, then a maximum Nash welfare allocation is EF1111 and PO and can be computed in polynomial time for binary additive valuations [DS15maximizing, BKV18greedy]. By contrast, a completion that maximizes Nash welfare could fail to be EF1111 even when an EF1111+PO completion exists (see Proposition 7 in Appendix A). Furthermore, the EF1111+PO-Completion problem turns out to be NP-complete. This follows from the construction of Theorem 1 where each item is allocated to an agent who approves it.
@@
Although Theorem 2 provides an efficient algorithm for checking the existence of a Prop1111 or an MMS completion, such a completion may not always exist (Proposition 8 in Appendix A). Our next result identifies a sufficient condition for guaranteeing the existence of an MMS+PO completion.
@@ 4 Lexicographic Valuations
Under lexicographic valuations, an allocation is Pareto optimal if and only if it satisfies sequencibility [HSV+21fair]. [^2] An allocation is said to be sequencible if there exists a picking sequence that induces it. Formally, an allocation $A$ is sequencible if there exists a string $\langle a_{1},a_{2},\dots,a_{m}\rangle$, where each $a_{i}\in N$ denotes the index of an agent, such that starting from $a_{1}$, if each agent $a_{i}$ picks its favorite remaining item on its turn, then the resulting allocation is $A$. Therefore, finding a PO allocation that completes a given frozen allocation boils down to finding a completion such that the final allocation has a corresponding picking sequence of agents. This can be done efficiently by constructing a picking sequence for the frozen allocation and extending it for the unallocated goods; see Theorem 10 in Appendix B. The following is an immediate consequence of this observation and Proposition 2.
@@
The detailed proof of Proposition 4 is deferred to Appendix B. Proposition 4 implies that MMS-Completion belongs to NP when agents have lexicographic valuations. Moreover, by exploiting the structure of an MMS partition, we establish the polynomial-time solvability of MMS-Completion. A crucial property here is that to ensure MMS for each agent, it is sufficient to allocate either the empty bundle, a single good, or a segment of their bottom-preferred unallocated goods. This implies that if an MMS allocation exists, there is one with at most one agent of the last type.
@@
To establish Theorem 5, we prove the following lemmas.
Lemma 1.
For an instance $\langle N,M,\mathcal{V},F\rangle$ with lexicographic valuations and agent $i\in N$, let $g_{1},g_{2},\ldots,g_{\ell}$ be the unallocated goods in $U=M\setminus\cup_{i\in N}F_{i}$ with $v_{i}(g_{1})>v_{i}(g_{2})>\ldots>v_{i}(g_{\ell})$. For every $C\subseteq U$ satisfying $v_{i}(F_{i}\cup C)\geq\mu_{i}$, either $C=\emptyset$, or there exists a good $g_{t}\in C$ such that
  • $v_{i}(F_{i}\cup\{g_{t}\})\geq\mu_{i}$, or
  • $\{g_{t},g_{t+1},\ldots,g_{\ell}\}\subseteq C$ with $v_{i}(F_{i}\cup\{g_{t},g_{t+1},\ldots,g_{\ell}\})=\mu_{i}$.
Proof of Lemma 1.
Lemma 2.
  • $(F_{1}\cup C_{1},\ldots,F_{n}\cup C_{n})$ is an MMS allocation,
  • |$|\{\,i\in N\mid|C_{i}|\geq 2\,\}|\leq 1$, and
Proof of Lemma 2.
We will now prove Theorem 5.
Proof of Theorem 5.
@@ 5 General Additive Valuations
Building on Proposition 5, we can show that for two agents with identical valuations, there is a polynomial-time algorithm for checking the existence of an EF1111 or Prop1111 completion. Note that EF1111 implies Prop1111 but the opposite may not hold. Indeed, an allocation that assigns two goods to agent $1$ and nothing to agent $2$ is Prop1111 but not EF1111. Nevertheless, a technique similar to Theorem 8 also works for Prop1111 and two agents with identical additive valuations.
@@
Proof.
$v(C_{1}\cup F_{1})$ $+\max_{j\in F_{2}}v(\{j\})$
$=v(C_{1})+v(F_{1})+\max_{j\in F_{2}}v(\{j\})$
$\geq\frac{1}{2}\left(v(C_{1})+v(C_{2})+v(F_{1})+v(F_{2})\right),$
As shown in Theorems 6 and 7, the EF1111-Completion problem is NP-complete for three agents with identical valuations or two agents with possibly non-identical valuations. Thus, our results present a tight demarcation of the frontier of tractability for the EF1111-Completion problem. Further, we note that EF1111-Completion problem for a general number of agents is strongly NP-complete; this can be observed by carrying out a reduction from 3-Partition.
@@
For additive valuations, an MMS allocation can fail to exist [KP18fair], but a $3/4$-MMS allocation always exists [AG24breaking]. By contrast, in the completion problem, it is impossible to guarantee $\alpha$-MMS for any $\alpha\in(0,1]$. Intuitively, the failure occurs for an instance where each agent $i$ thinks that the first $i$ agents should equally divide the unallocated goods and the remaining agents are already rich enough. This discrepancy in agents’ valuations prohibits the existence of an $\alpha$-MMS completion.
@@
In the standard setting of fair division, the decision problem of determining the existence of an MMS allocation is in $\Sigma^{2}_{p}$ [BL16characterizing]; however, it remains open whether the problem is NP-hard or even harder. A simple reduction from PARTITION shows that the search version of the problem is NP-hard. It is easy to see that MMS-Completion includes the above decision problem as a special case. Our final result shows that MMS-Completion is NP-hard even for two agents with additive valuations. The proof of this result is deferred to Appendix C.
@@ Acknowledgments
We thank anonymous reviewers for helpful comments. Ayumi Igarashi acknowledges support from JST FOREST under grant number JPMJFR226O. Rohit Vaish acknowledges support from DST INSPIRE grant no. DST/INSPIRE/04/2020/000107 and SERB grant no. CRG/2022/002621.
@@ Appendix B Omitted Material from Section 4
    $a_{e}:\{v_{j_{1}},\dots,v_{j_{\ell}}\}\succ\langle\text{all frozen hyperedge and all frozen pair goods except $g_{e}$}\rangle\succ\underline{g_{e}}\succ\langle\text{rest}\rangle.$
@@
    $p_{e,v}:v\succ g_{e}\succ\langle\text{all frozen pair goods except for }g_{e,v}\rangle\succ\underline{g_{e,v}}\succ\langle\text{rest}\rangle.$
@@ Appendix C Omitted Material from Section 5
$\mu_{i}\geq\epsilon+\mu^{*}.$

### v2

@@ Abstract
We formulate the problem of fair and efficient completion of indivisible goods, defined as follows: Given a partial allocation of indivisible goods among agents, does there exist an allocation of the remaining goods (i.e., a completion) that satisfies fairness and economic efficiency guarantees of interest? We study the computational complexity of the completion problem for prominent fairness and efficiency notions such as envy-freeness up to one good (EF1111), proportionality up to one good (Prop1111), maximin share (MMS), and Pareto optimality (PO), and focus on the class of additive valuations as well as its subclasses such as binary additive and lexicographic valuations. We find that while the completion problem is significantly harder than the standard fair division problem (wherein the initial partial allocation is empty), the consideration of restricted preferences facilitates positive algorithmic results for threshold-based fairness notions (Prop1111 and MMS). On the other hand, the completion problem remains computationally intractable for envy-based notions such as EF1111 and EF1111+PO even under restricted preferences.
@@ 1 Introduction
Another challenge posed by the completion problem is in the formulation of solution concepts. Specifically, let us consider the maximin share (MMS) criterion [17], where fairness is defined based on a utility threshold for each agent. This threshold is defined in terms of an $n$-person cut-and-choose game, where the cutter is the last to pick a bundle. When some resources are preassigned, there can be multiple ways of extending this definition. For example, the cutter may only consider the extension of its own frozen bundle, or it may seek to maximize the value of the least-valued bundle overall. These choices lead to different existential and computational results. Thus, adapting existing solution concepts to the completion setting is non-trivial.
@@
EF1111 Poly NP-c (Th. 1) Poly NP-c (Th. 3) Poly NP-c (Th. 6,7)
Prop1111 Poly Poly (Th. 2) Poly Poly (Th. 4) Poly NP-c (Th. 6,7)
MMS Poly§ Poly (Th. 2) Poly Poly (Th. 5) ${\Sigma^{2}_{p}}^{\S}$ NP-h (Th. 9)
EF1111+PO Poly NP-c (Cor. 1) Poly ? Pseudopoly NP-h (Cor. 3)
Prop1111+PO Poly Poly (Cor. 2) Poly Poly (Th. 4) Poly NP-h (Cor. 3)
MMS+PO Poly§ Poly (Cor. 2) Poly ? ? ?
@@
  • For general additive valuations, we show that the completion problem for EF1111 and Prop1111 (and their combinations with Pareto optimality) is NP-hard. This hardness result holds even for two agents with additive valuations and for three agents with identical additive valuations. We also establish the hardness of deciding the existence of an MMS allocation, a problem whose lower bound on complexity is open in the standard setting. Moreover, we demonstrate that $\alpha$-MMS may not exist for every $\alpha>0$ even when the initial allocation is Pareto optimal.
@@ 3 Binary Additive Valuations
In the standard setting of fair division (i.e., without any frozen goods), an EF1111 allocation always exists and can be efficiently computed by a simple round-robin algorithm [CKM+19unreasonable]. However, as discussed in the Introduction, an EF1111 allocation could fail to exist when some goods are pre-allocated. Our first main result in this section is that EF1111-Completion is computationally intractable even when the agents have binary additive valuations and even when the frozen allocation is Pareto optimal. Note that, for binary additive valuations, an allocation is Pareto optimal if and only if each good is allocated to an agent who approves it.
@@
If there is no frozen good, then a maximum Nash welfare allocation is EF1111 and PO. For binary additive valuations, a maximum Nash welfare allocation can be computed in polynomial time [DS15maximizing, BKV18greedy]. By contrast, a completion that maximizes Nash welfare could fail to be EF1111 even when an EF1111+PO completion exists; see Proposition 7 in Appendix A. Furthermore, the EF1111+PO-Completion problem turns out to be NP-complete. This follows from the construction of Theorem 1 where each item is allocated to an agent who approves it.
@@
Although Theorem 2 provides an efficient algorithm for checking the existence of a Prop1111 or an MMS completion, such a completion may not always exist; see Proposition 8 in Appendix A. Our next result identifies a sufficient condition for guaranteeing the existence of an MMS+PO completion.
@@ 4 Lexicographic Valuations
The proof of Theorem 3 uses ordinal valuations. Thus, if we do not require cardinal valuations in the definition of lexicographic preferences, EF1111-Completion becomes strongly NP-hard. However, if we represent the valuations as numbers (for example, powers of 2), it is not clear whether strong NP-hardness holds for lexicographic valuations.
@@
Under lexicographic valuations, an allocation is Pareto optimal if and only if it satisfies sequencibility [HSV+21fair]. [^2] An allocation is said to be sequencible if there exists a picking sequence that induces it. Formally, an allocation $A$ is sequencible if there exists a string $\langle a_{1},a_{2},\dots,a_{m}\rangle$, where each $a_{i}\in N$ denotes the index of an agent, such that starting from $a_{1}$, if each agent $a_{i}$ picks its favorite remaining item on its turn, then the resulting allocation is $A$. Therefore, the problem of finding a PO allocation that completes a given frozen allocation reduces to finding a completion such that the final allocation has a corresponding picking sequence of agents. The latter problem can be solved efficiently by constructing a picking sequence for the frozen allocation and extending it for the unallocated goods (see Theorem 10 in Appendix B). The following is an immediate consequence of this observation and Proposition 2.
@@
The detailed proof of the above proposition is deferred to Appendix B. Proposition 4 implies that MMS-Completion belongs to NP when agents have lexicographic valuations. Moreover, by exploiting the structure of an MMS partition, we establish the polynomial-time solvability of MMS-Completion. A crucial property here is that to ensure MMS for each agent, it is sufficient to allocate either the empty bundle, a single good, or a segment of their bottom-preferred unallocated goods. This implies that if an MMS allocation exists, there is one with at most one agent of the last type.
@@ 5 General Additive Valuations
Note that the two assumptions in Proposition 5 are necessary to guarantee the existence of an EF1 completion and its polynomial-time computability. In fact, if the frozen allocation is EF1111 but the envy graph contains a cycle, the completion problem becomes computationally hard. For example, we can construct an instance where the frozen allocation is EF1111, but every agent envies another (e.g., each agent receives an item valued at 0, while all other agents receive items valued at a very large amount from their perspective). This setup essentially reduces the problem to deciding whether there exists an envy-free allocation of the unallocated items, which is known to be NP-hard even for binary additive valuations [BouveretL08]. Moreover, if the envy graph of the frozen allocation is acyclic, our Theorem 7 already demonstrates that the completion problem remains hard (recall that for identical valuations, any allocation is Pareto-optimal).
Building on Proposition 5, we can show that for two agents with identical valuations, there is a polynomial-time algorithm for checking the existence of an EF1111 or Prop1111 completion. Note that EF1111 implies Prop1111 but the opposite may not hold. Indeed, an allocation that assigns two goods to agent $1$ and nothing to agent $2$ is Prop1111 but not EF1111. Nevertheless, a technique similar to that for EF1111 also works for Prop1111 and two agents with identical additive valuations.
@@
As shown in Theorems 6 and 7, the EF1111-Completion problem is NP-complete for three agents with identical valuations or two agents with possibly non-identical valuations. Thus, our results present a tight demarcation of the frontier of tractability for the EF1111-Completion problem. Further, we note that EF1111-Completion problem for a general number of agents is strongly NP-complete; this can be observed by carrying out a reduction from 3-Partition. This result is presented as Theorem 11 in Appendix C.
@@
For additive valuations, an MMS allocation can fail to exist [KP18fair], but a $(3/4+3/3836)$-MMS allocation always exists [AG24breaking]. By contrast, in the completion problem, it is impossible to guarantee $\alpha$-MMS for any $\alpha\in(0,1]$. Intuitively, the failure occurs for an instance where each agent $i$ thinks that the first $i$ agents should equally divide the unallocated goods and the remaining agents are already rich enough. This discrepancy in agents’ valuations prohibits the existence of an $\alpha$-MMS completion.
@@
In the standard setting of fair division, the decision problem of determining the existence of an MMS allocation is in $\Sigma^{2}_{p}$ [BL16characterizing]; however, it remains open whether the problem is NP-hard or even harder. A simple reduction from PARTITION shows that the search version of the problem is NP-hard. It is easy to see that MMS-Completion includes the above decision problem as a special case. Our final result shows that MMS-Completion is NP-hard even for two agents with additive valuations.
@@ 6 Conclusion
In our paper, we defined the notion of MMS, which incorporates a pre-determined frozen allocation. An alternative approach would be to use the standard MMS definition that disregards the frozen allocation and study the problem of achieving this MMS. Notably, with the standard MMS definition, our computational complexity results remain unchanged. For binary additive valuations, it is still possible to compute an MMS threshold in polynomial time, allowing us to apply the same network-flow argument to compute an MMS allocation. For lexicographic valuations, under the MMS definition, each agent’s MMS partition should take the form ($(\{g_{1}\},\{g_{2}\},\{g_{n},\ldots,g_{m}\})$ [HSV+20fairAAAI] where $g_{1}$ is the most favorite good, and $g_{2}$ is the second favorite good, and so on. After removing agents whose frozen bundles already achieve their MMS threshold, one can iteratively guess which agent should receive a bundle of $m-n+1$ items, create a bipartite graph where there is an edge between an unallocated item $g$ and an agent $i$ such that $i$ has an MMS value for $g$, and decide whether there exists a matching that saturates the agents in the graph. For general additive valuations, the NP-hardness of deciding the existence of an MMS allocation (Theorem 9) still holds. This is because each agent assigns a very small value to the frozen items, resulting in the same thresholds for both our MMS definition and the standard MMS definition.
@@ Acknowledgments
We thank anonymous reviewers for helpful comments. Ayumi Igarashi acknowledges support from JST FOREST under grant number JPMJFR226O. Rohit Vaish acknowledges support from DST INSPIRE grant no. DST/INSPIRE/04/2020/000107 and SERB grant no. CRG/2022/002621. Vishwa Prakash HV acknowledges support of TCS Research Scholar Fellowship.
@@ Appendix B Omitted Material from Section 4
    $a_{e}:\{v_{j_{1}},\dots,v_{j_{\ell}}\}\succ$
[…]
    $\succ\underline{g_{e}}\succ\langle\text{rest}\rangle.$
@@
    $p_{e,v}:v\succ g_{e}$ $\succ\langle\text{all frozen pair goods except for }g_{e,v}\rangle$
    $\succ\underline{g_{e,v}}\succ\langle\text{rest}\rangle.$
@@
$\qquad v_{i}((F_{j^{\prime\prime}}\cup C_{1}\cup C_{j^{\prime\prime}})\setminus\{g_{1}\}),\min_{j\neq 1,j^{\prime\prime}}v_{i}(F_{j}\cup C_{j})\}$
@@
To establish Theorem 5, we prove the following lemmas.
Lemma 3.
For an instance $\langle N,M,\mathcal{V},F\rangle$ with lexicographic valuations and agent $i\in N$, let $g_{1},g_{2},\ldots,g_{\ell}$ be the unallocated goods in $U=M\setminus\cup_{i\in N}F_{i}$ with $v_{i}(g_{1})>v_{i}(g_{2})>\ldots>v_{i}(g_{\ell})$. For every $C\subseteq U$ satisfying $v_{i}(F_{i}\cup C)\geq\mu_{i}$, either $C=\emptyset$, or there exists a good $g_{t}\in C$ with
  • $v_{i}(F_{i}\cup\{g_{t}\})\geq\mu_{i}$, or
  • $\{g_{t},g_{t+1},\ldots,g_{\ell}\}\subseteq C$ such that $v_{i}(F_{i}\cup\{g_{t},g_{t+1},\ldots,g_{\ell}\})=\mu_{i}$.
Proof.
Lemma 4.
  • $(F_{1}\cup C_{1},\ldots,F_{n}\cup C_{n})$ is an MMS allocation,
  • |$|\{\,i\in N\mid|C_{i}|\geq 2\,\}|\leq 1$, and
Proof.
We will now prove Theorem 5.
Proof of Theorem 5.
@@ Appendix C Omitted Material from Section 5
$v(C_{1}\cup F_{1})$ $+\max_{j\in F_{2}}v(\{j\})$
$=v(C_{1})+v(F_{1})+\max_{j\in F_{2}}v(\{j\})$
$\geq\frac{1}{2}\left(v(C_{1})+v(C_{2})+v(F_{1})+v(F_{2})\right),$
Theorem 11.
For a general number of agents with identical additive valuations, EF1111-Completion and Prop1111-Completion are strongly NP-complete.
Proof.
EF1111: The problem is clearly in NP. To show strong NP-hardness, we present a reduction from 3-Partition. An instance of 3-Partition is given by a set of $3n$ numbers $w_{1},w_{2},\dots,w_{3r}$ where $r\in\mathbb{N}$. Let $T$ be such that $\sum_{i\in[3r]}w_{i}=rT$. The goal is to determine if there exists a partition of $w_{1},w_{2},\dots,w_{3r}$ into $r$ triples such that the sum of numbers in each triple is $T$. This problem is known to be NP-complete even when, for every $i\in[3r]$, we have $w_{i}\approx T/3$. Specifically, there is a fixed constant $\delta>0$ such that for every $i\in[3r]$, we have $w_{i}\in[T/3-\delta,T/3+\delta]$, see [strong3partition].
Given any instance of 3-Partition, we will create an instance of the completion problem as follows. There are $r+1$ agents, namely $11,2,\dots,r+$ 1, with the same valuation function $v$. There are two frozen goods $f_{1}$ and $f_{2}$, each valued at $T$ by all agents, that are allocated to agent $r+1$. For every $k\in[3r]$, we create a good $g_{k}$ with the value $w_{k}$ for all agents.
We will now argue that the given 3-Partition instance admits the desired partition into triples if and only if an EF1111 completion exists in the fair division instance. For the proof of forward direction, let’s assume that there exists a partition into triples $S_{1},S_{2},\dots,S_{r}$ such that the sum of numbers in each triple is $T$. For each triple $S_{i}$, we will assign the goods corresponding to the numbers in $S_{i}$ to agent $i$. Thus, for every $i\in[r]$, agent $i$’s value for its bundle is $T$. Additionally, for any $k\neq r+1$, agent $i$ values agent $k$’s bundle is also $T$, while its value for agent $(r+1)$’s bundle after removing the frozen good $f_{1}$ is $T$. It is easy to see that agent $(r+1)$ does not envy any other agent. Thus, the completion is EF1111.
In the reverse direction, suppose an EF1111 completion exists. Let us denote the final allocation by $A=(F_{i}\cup C_{i})_{i\in[r+1]}$. Since $A$ is EF1111 and the valuations are identical, each agent in $\{1,\dots,r\}$ must receive a value of at least $T$, as otherwise such agents would envy agent $r+1$ even after removal of any good. The total value of the unfrozen goods is $rT$, which means that all unfrozen goods must be allocated to the agents in $\{1,\dots,r\}$. Furthermore, each agent in $\{1,\dots,r\}$ must receive a value of exactly $T$. To obtain a value of $T$, each agent must receive at least three goods (recall that each good’s value is nearly $T/3$). This assignment of goods naturally induces a valid partition into triples of the integers $w_{1},\dots,w_{3r}$.
Prop1111: To show strong NP-hardness of Prop1111-Completion, we will again present a reduction from 3-Partition. We will assume that the numbers $w_{1},\dots,w_{3r}$ in 3-Partition are such that for every $i\in[3r]$, we have $w_{i}>>1$; this property can be ensured by appropriate scaling.
Given an instance $w_{1},\dots,w_{3r}$ of 3-Partition, we will create a fair division instance with $r+1$ agents, namely $11,2,\dots,r+$ 1, with identical valuations. There are $1+(r+1)T$ frozen goods, all of them allocated to agent $r+1$, consisting of one good of value $T$ and $(r+1)T$ other goods each of value $1$. Additionally, there are $3r$ unfrozen goods corresponding to the $3r$ integers $w_{1},\dots,w_{3r}$. The $i^{\textup{th}}$ such good is valued at $w_{i}$.
Let us first present the argument for the forward direction. Note that the proportionality threshold in the fair division instance is $\theta\coloneqq\frac{rT+T+(r+1)T}{r+1}=2T$. Given a desired partition into triples $S_{1},\dots,S_{r}$ in the 3-Partition instance, we create a Prop1111 allocation as follows: Assign the goods corresponding to the numbers in $S_{i}$ to agent $i$. This gives each agent in $\{1,\dots,r\}$ a utility of $T$. Each such agent can reach the proportionality threshold $\theta=2T$ by hypothetically including the frozen good valued at $T$ owned by agent $(r+1)$. Since agent $(r+1)$ already satisfies proportionality, the resulting allocation is Prop1111.
In the reverse direction, suppose a Prop1111 allocation exists. Let us denote this allocation by $A=(F_{i}\cup C_{i})_{i\in[r+1]}$. We claim that each agent in $\{1,\dots,r\}$ must derive a utility of at least $T$ under $A$. This is because the proportionality threshold is $\theta=2T$, and the largest good available to any agent in $\{1,\dots,r\}$ for hypothetical addition under Prop1111 is of value $T$ (this is the frozen good owned by agent $r+1$). By a similar argument as in the EF1111 reduction above, it can be observed that any such allocation induces the desired partition into triples in the given instance of 3-Partition. ∎